OpenStudy (anonymous):

The area of a rectangle must be at least 153 square feet. If the length is 9 feet, find the minimum for the rectangle's width.

OpenStudy (anonymous):

Can anyone help

OpenStudy (jdoe0001):

 $$ \text{since area is }\color{blue}{length \times width}\\ 9\times x \ge 153\\ \text{solve for "x"} $$

OpenStudy (anonymous):

And I'll get my answer?

OpenStudy (jdoe0001):

yes, is really just a plain area formula issue, the only thing is worded "at least 153" so if something is at least 153, it can't be 152, but it can be 153 or 154 or 2000 so it can really be greater than or equal to that